OpenStudy (anonymous):

Solve the equation. 7(x + 3) = (7x + 21)

OpenStudy (anonymous):

\[7(x+3)=7(x)+7(3)\] \[=?\]

OpenStudy (anonymous):

distribute 7 on the left side 7x+21=7x+21 some typo is there . check the question again.

OpenStudy (anonymous):

is it 7(x+3)=0 ??

OpenStudy (anonymous):

0?? i think

OpenStudy (anonymous):

o just seen your reply i got 0 too thanks

OpenStudy (anonymous):

what the :D it's already done

OpenStudy (anonymous):

Is valid for all values of x.

OpenStudy (anonymous):

these are my choises A. 0 B. 42 C. All real numbers D. No solution

OpenStudy (anonymous):

A correct?

OpenStudy (anonymous):

no i think E is correct

OpenStudy (anonymous):

It's C.

OpenStudy (anonymous):

, that equation obeys. for any x-value you put in there

OpenStudy (anonymous):

No matter what x-value you put in the equation you were given, it will equal each other. That means the equation is true.

OpenStudy (anonymous):

oo ok i see now -_- not much of a math person thanks for the help